# An Empirical Study of Self-Admitted Technical Debt in Machine Learning Software

### v2

@@ 4.5.1. RQ1: What is the prevalence of SATD in ML-based systems?
We extract 318 ML and 318 non-ML python projects following the description from Section 4. For each ML or non-ML project, we extract comments using Comment-Parser, a PyPi framework [^6] Accessed on: 2024-03-22, https://pypi.org/project/comment-parser/. We obtained 0.4 Million and 1.5 Million comments in ML and non-ML software respectively. A detailed description of the distribution of ML and non-ML software parameters is provided in Table 3. We qualify whether a comment is SATD using the SATD detection tool [^7] Accessed on: 2024-03-22, https://github.com/Tbabm/SATDDetector-Core. From Table 3, we observe that the number of comments in non-ML projects is approximately 2.8 times that of ML projects. To resolve the imbalance in this comparison, we collect a sample of 100,000 ML and non-ML comments and compared the percentage of SATD in ML and non-ML projects. We repeat this 100 times, to ensure that our results are not based on a particular sample.
@@
A common approach to challenge prevailing assumptions and explore alternative explanations for an empirical observation is the Advocatus Diaboli (a.k.a. the devil’s advocate) approach (vysali2020quantifying, ). This approach involves identifying the simplest, most logical explanations for a given observation, then hypothesizing and validating whether these might explain it. In our case, we adopt the devil’s advocate perspective to examine the high debt prevalence in ML and Non-ML settings by assessing three potential predictors: 1) churn, 2) number of file changes, and 3) complexity. While other metrics could be relevant, our selection is designed to provide an evaluation of these particular variables, ensuring a focused and manageable scope for detailed investigation. Below, we present the rationale to select our metrics in the context of technical debt.
  • Churn. High churn indicates frequent changes and adjustments, which we hypothesize could potentially lead to a breeding ground for technical debt. High churn is associated with other poor programming practices and design issues, including higher likelihood of defects and smells (amit2020corrective, ; kudrjavets2022quantifying, ; mauerer2021search, ; Morales15, ). Employing the broken window theory (leven2022broken, ), high churn signifying other aspects of poor design could also be extended to technical debt.
@@
  • Complexity of File. Complexity directly hampers the understandability and thus the maintainability of code. Research has indicated that quality attributes like debugging effort, security, vulnerability, refactoring effort, fault density, fault proneness, and maintainability are impacted by complexity (nguyen2017impact, ). Hence, high complexity code is hypothesized to be prone to accruing technical debt.
@@ 4.5.2. RQ2: What are the different types of SATD in ML-based systems?
They used card sorting to categorize the SATDs using the extended taxonomy proposed by Bavota and Russo (bavota2016large, ). The latter taxonomy includes SATD types by Bavota along with the addition of 1) “On-Hold Debt” (Rungroj, ), and 2) “Refactoring” under “Code Debt” (suryanarayana2014refactoring, ) (see Figure 9).
In cases where the SATDs could not be categorized using the labels proposed by this taxonomy, a new category was created by us. A similar process was adopted by Bhatia et al. (bhatia2023towards, ) for the expansion of a prior taxonomy of code changes by Hindle et al. (hindle2008large, ) to involve the ML types of code changes. Akin to our research, Bhatia et al. (bhatia2023towards, ) employed card sorting, introducing new types of changes as additional codes whenever the existing labels did not adequately describe the changes.
After this initial step, the authors separately classified another set of 65 samples (from a pool of 611), then computed the inter-rater agreement on this new sample using Cohen Kappa, and obtained a value of 94%, which is very high. After achieving this high level of agreement, the first author manually classified the remaining 546 SATD comments. A similar approach was used by Kononenko et al. (7886977, ) to classify code review survey responses, and Uddin and Khomh (8643972, ) to classify API aspects.
@@ 4.5.3. RQ3: Which stages of the ML pipeline are more prone to SATD?
We utilized the same dataset of 611 SATD comments generated for RQ2, with the objective of mapping these comments to specific stages of the ML workflow. Following the categorization established by Amershi et al. (amershi2019software, ), we leverage five critical stages within the ML lifecycle, namely: data reading, data preprocessing, model building, model validation, and model deployment. A similar approach was followed by Foalem et al. (foalem2024studying, ) for mapping log statements to a corresponding ML pipeline stage.
@@
To ensure reliability in our categorization, the first two authors conducted an initial pilot study on an initial set of 65 samples to independently provide labels, and later discussed and resolved any classification disputes until consensus was reached. The inter-rater agreement, measured by Cohen’s Kappa, was 81% among the first two authors for a subset of 65 comments. This level of agreement, surpassing 70%, is considered adequate (byrt1996good, ). Following this pilot phase, we refined our categorizations by 1) eliminating any false positives, 2) removing the SATD instances labeled “unknown” and 3) removing the SATD from Non-ML components. Eventually, we found a total of 370 samples corresponding to different stages of the ML pipeline.
@@ 4.5.4. RQ4: How long does SATD survive in ML-based systems?
We perform survival analysis (miller2011survival, ) to examine, for each project, (i) the time until the introduction of SATD and (ii) the time taken to remove SATDs after they are introduced in a file for the first time. The former analysis examines the survival of ML software projects concerning debt occurrence, while the latter analysis is about the survival of SATDs instances once they are introduced in an ML-based software project.
Survival analysis (miller2011survival, ) is a statistical technique for analyzing the expected duration of time until the occurrence of an event of interest or censoring of an event (if the event does not happen). The event of interest in our study is the occurrence of SATD across the entire project history, and the length of the observation window is the start of the project until the time at which we obtained the projects’ latest snapshot at the time of performing the empirical study (Summer 2021). If the event is not observed during the observation period, the corresponding subject will be censored at the end of the period, which survival analysis is able to deal with.
@@
Time to event and status are two important variables for survival analysis. Time to event ($T$) is the elapsed time between the beginning of the observation (start date of the project) and the occurrence of the event of interest or the censoring of data. In our study, this time is measured in #days. Hence, $T$ is a random variable with positive values (miller2011survival, ). A summary of the Time ($T$), Event of interest, and Status is provided in Table 4. For the introduction of SATD, $T$ is the elapsed time between the start date of the project to the time of contamination of the file by SATD. For the removal of SATD, $T$ is the elapsed time between the removal and the introduction of SATD.
@@
The Survival function S(t) gives the probability ($P(T>t)$) that a subject will survive beyond time $t$. After arranging our data in increasing order of $T$, we plot the survival curve and estimate the survival probability using the non-parametric Kaplan-Meier estimator (kaplan1958nonparametric, ). The Kaplan-Meier estimation is computed following Equation 1, where $t_{i}$ is the time duration up to event-occurrence point $i$, $d_{i}$ is the number of event occurrences up to $t_{i}$, and $n_{i}$ is the number of subjects that survive just before $t_{i}$. $n_{i}$ and $d_{i}$ are obtained from the aforementioned ordered data.
@@ 4.5.5. RQ5: What are the characteristics of long-lasting SATDs in ML-based systems?
Our dimensions ‘Change’, ‘Diffusion’, ‘Complexity’, and ‘History’, are widely employed in the domain of SDP (software defect prediction), a well researched field in Software Engineering. The rationale for their usage in the context of explaining SATD survival is based on the understanding that areas of code that undergo frequent changes, or are complex, are more prone to defects. Given that SATD indicates sub-optimal implementations that may not be outright defects but could result in compromised software quality, we consider using these metrics for SATD prediction. Moreover, prior research in SATD prediction by Yan et al. (yan2018automating, ) previously employed such dimensions (i.e., diffusion, history, and message) in their study. In contrast to Yan et al.’s research, we exclude the message dimension from our study as our focus is on assessing the attributes of persistent self-admitted technical debt (SATD) rather than SATD detection itself.
@@
Assigning Class Labels
To identify long-lasting SATDs, we compute the time elapsed since the creation of the comment admitting the debt. We compute this elapsed time using information from the commit diffs (i.e., +++++ + and --- - lines of code comments for each source code file in the commit). Specifically, we calculate the time between the introduction of the SATD comment and its removal. From our collected data, we observed cases in which the same SATD commit is present multiple times within the same file, and removed these cases from our case study to avoid biasing our results with duplicated information. After computing SATD duration time for a total of 6,427 SATD instances in our dataset, we computed the quartiles of the obtained distribution and used the first and fourth quartile to identify “quick removal” and “long lasting” SATD respectively. In total, we obtained 2,595 SATD instances with duration times above 75th percentile. We assigned them to the group of long-lasting SATDs. We obtained the first quantile, i.e., 2,595 instances of with duration times below the 25th percentile as the quick removal SATD.
Correlation and Redundancy Analysis
@@
Model Validation
To ensure the statistical robustness of our results, we perform a 100-out-of-sample bootstrapping. The approach involves sampling the training set from the whole dataset with replacement. The testing set comprises all the samples present in the original data but absent in the bootstrapped training data. We build and test models on each training and testing set to draw statistically robust conclusions as suggested by (tantithamthavorn2016empirical, )
Model Interpretation
We use SHAP (SHapely Additive exPlanations), to interpret and explain our models. A Shapley value is the average marginal contribution of a feature value across all possible coalitions (shapley201617, ). We compute these values using the Python framework, SHAP [^14] Accessed on: 2024-03-22, https://github.com/slundberg/shap which provides a collaborative game theory approach for the calculation of the average marginal contribution of each of the model’s features.
@@ 4.6. Replication Package
All the scripts along with the mined data is provided in the replication package [^15] Accessed on: 2024-03-22, https://drive.google.com/drive/folders/1n-gwAxFANS-PPewnk1Qgmnwd0zMAr7Yb *This tentative g-drive link will be replaced by a github repository after the reviewing process.
@@ 5.1. RQ1: What is the prevalence of SATD in ML-based systems?
Figure 5. Percentage of SATD in ML and non-ML projects. The comparison has been made by collecting 100,000 random samples from ML and non-ML comments and repeating the process 100 times (to ensure robustness).
Machine learning-based projects have 2.1 times more SATD in comparison to non machine learning projects.
@@
Figure 5 shows the distribution of SATD in ML and non-ML software. The median of percentages of SATD in ML and non-ML software is 1.9% and 0.9%, calculated over 100 random samples of 100,000 comments from ML and non-ML respectively. This indicates a 2.1 times more debt in ML projects. The comparisons provided in Figure 5 are statistically significant with the Wilcoxon Signed Rank test $p$-value lower than 0.05 and a large Cohen’s $d$ effect size difference, but did not impact the ranges for $p$-value or the effect size. Changing the sample size of 100,000 did not change our results of median values, but only changed the standard deviation of both distributions in Figure 5. When considering all the comments in ML and non-ML code from our dataset (without any sampling), 3.5% of comment declarations in ML projects contain a technical debt admission (i.e., SATD), compared to only 2% for non-ML projects. Our findings for SATD in non-ML software are aligned with the findings of prior research. For instance, Bavota and Russo (bavota2016large, ) found SATD in 0.3% of the comments of non-ML Java-based projects, which is comparable to the 0.9% obtained by our analysis for Python-based non-ML projects.
Since ML developers tend to experiment with different configurations when searching for the optimal model, we surmise that the amount of code churn in ML code is likely higher than the amount of code churn in non-ML code, which could explain the observed difference in the proportion of SATD (since high code churn is known to be associated with poor design ). Hence, we computed the distribution of code churn in the studied ML and non-ML projects.
The results presented in Figure 6(a) show that ML code experiences statistically significantly higher code churn than non-ML code, with large effect size (Cohen’s $\delta$=12). One possible explanation of the high rate of SATD in ML code could be the higher rate of code modification, even though we cannot claim any causal relation. Furthermore, Figure 6 (b) shows that SATD files also have higher churn-per-file size for ML software. This further corroborates that those larger changes correlate with lower software quality and higher amounts of SATD. The comparisons of churn and churn-per-LOC in Figure 6 (a and b) are statistically significant with large effect size for SATD code and medium effect size for clean code.
Additionally, we looked into alternative explanations of the higher amounts of SATD in ML by checking alternative metrics like file-level complexity and commit-level file changes. While both code complexity and the number of file changes per commit are lower for ML than for non-ML code, MLcode churn is significantly higher than non-ML churn. Figure 7 (a) shows the normalized number of files changed in a commit, while Figure 7 (b) shows the normalized complexity of the changed ML and non-ML files. Although the comparison between ML and non-ML file counts are significant, there is only a negligible effect size. In Figure 7 (b), when comparing ML to non-ML code complexity using Cohen’s $\delta$, the clean files, represented by the red boxes, show a large effect size with a value of 0.8. On the other hand, the SATD files, represented by the blue boxes, show a small effect size with a value of 0.2.
Overall, this indicates that ML and non-ML software projects differ along metrics like complexity and churn, and ML churn is significantly higher than non-ML churn despite the ML complexity being significantly lower than non-ML. The churn in machine learning (ML) development could be notably higher than in non-ML code largely due to the repetitive nature of ML-specific tasks such as data preprocessing, model building, and hyperparameter tuning. These activities often necessitate frequent code modifications, as highlighted by Sculley et al. (Sculley, ), with even minor conceptual changes leading to substantial code alterations. This is evident in scenarios like dataset modifications, where changes to a feature column can require extensive updates to associated code, or during model building, where continuous adjustments to hyperparameters contribute to churn. Moreover, switching between ML frameworks, like from Scikit-learn to TensorFlow, exacerbates this issue due to significant differences in their syntax and functions. This high churn rate not only parallels the negative impact of code churn on traditional software quality, as shown in various studies (mauerer2021search, ; Morales15, ), but can also affect ML software by potentially degrading quality of ML-based systems. ML developers should be mindful of the deterioration of code quality when making frequent changes to their codebase. They should consider monitoring their technical debt, to prevent its cost from reaching a critical level quickly.

### v3

@@ 4.5.1. RQ1: What is the prevalence of SATD in ML-based systems?
We extract 318 ML and 318 non-ML python projects following the description from Section 4. For each ML or non-ML project, we extract comments using Comment-Parser, a PyPi framework [^8] Accessed on: 2024-03-22, https://pypi.org/project/comment-parser/. We obtained 0.4 Million and 1.5 Million comments in ML and non-ML software respectively. A detailed description of the distribution of ML and non-ML software parameters is provided in Table 3. We qualify whether a comment is SATD using the SATD detection tool [^9] Accessed on: 2024-03-22, https://github.com/Tbabm/SATDDetector-Core. We perform sampling with replacement (bootstrapping) and compare the amount of SATD in ML and Non-ML projects across 1,000 runs. Since the number of comments in ML and Non-ML projects are disproportionate (see Table 3), bootstrapping prevents our results to be dependent on any one-specific sample seed, keeping the Standard Error ($\mathrm{SE}=\frac{s}{\sqrt{n}}$) of our setup as low as 0.07% for ML comments and 0.05% for non-ML comments.
@@
A common approach to challenge prevailing assumptions and explore alternative explanations for an empirical observation is the Advocatus Diaboli (a.k.a. the devil’s advocate) approach (vysali2020quantifying). This approach is a recognized methodology used in prior software engineering and broader scientific research to systematically test alternative hypotheses. For instance, Minelli et al. (minelli2015know) used it to question whether their findings generalized beyond a Smalltalk IDE; Ferré and Rudolph (ferre2012advocatus) applied it to falsify the completeness of ontologies via logically absurd counterexamples. In line with these works, we apply this technique not to claim causation via metrics like churn or complexity, but to challenge the robustness of our own explanation by actively testing plausible alternatives.
In our case, we adopt the devil’s advocate perspective to examine the high debt prevalence in ML and Non-ML settings by assessing three potential predictors: 1) churn, 2) number of file changes, and 3) complexity. While other metrics could be relevant, our selection is designed to provide an evaluation of these particular variables, ensuring a focused and manageable scope for detailed investigation. Below, we present the rationale to select our metrics in the context of technical debt.
  • Churn. High churn indicates frequent changes and adjustments, which we hypothesize could potentially lead to a breeding ground for technical debt. High churn is associated with other poor programming practices and design issues, including higher likelihood of defects and smells (amit2020corrective; mauerer2021search; Morales15). Employing the broken window theory (leven2024broken), high churn signifying other aspects of poor design could also be extended to technical debt.
@@
  • Complexity of File. Complexity directly hampers the understandability and thus the maintainability of code. Research has indicated that quality attributes like debugging effort, security, vulnerability, refactoring effort, fault density, fault proneness, and maintainability are impacted by complexity (nguyen2017impact). Hence, high complexity code is hypothesized to be prone to accruing technical debt.
@@ 4.5.2. RQ2: What are the different types of SATD in ML-based systems?
They used card sorting to categorize the SATDs using the extended taxonomy proposed by Bavota and Russo (bavota2016large). The latter taxonomy includes SATD types by Bavota along with the addition of 1) “On-Hold Debt” (Rungroj), and 2) “Refactoring” under “Code Debt” (suryanarayana2014refactoring) (see Figure 9).
In cases where the SATDs could not be categorized using the labels proposed by this taxonomy, a new category was created by us. A similar process was adopted by Bhatia et al. (bhatia2023towards) for the expansion of a prior taxonomy of code changes by Hindle et al. (hindle2008large) to involve the ML types of code changes. Akin to our research, Bhatia et al. (bhatia2023towards) employed card sorting, introducing new types of changes as additional codes whenever the existing labels did not adequately describe the changes.
After this initial step, the authors separately classified another set of 65 samples (from a pool of 611), then computed the inter-rater agreement on this new sample using Cohen Kappa, and obtained a value of 94%, which is very high. After achieving this high level of agreement, the first author manually classified the remaining 546 SATD comments. A similar approach was used by Kononenko et al. (7886977) to classify code review survey responses, and Uddin and Khomh (8643972) to classify API aspects.
@@ 4.5.3. RQ3: Which stages of the ML pipeline are more prone to SATD?
We utilized the same dataset of 611 SATD comments generated for RQ2, with the objective of mapping these comments to specific stages of the ML workflow. Following the categorization established by Amershi et al. (amershi2019software), we leverage five critical stages within the ML lifecycle, namely: data reading, data preprocessing, model building, model validation, and model deployment. A similar approach was followed by Foalem et al. (foalem2024studying) for mapping log statements to a corresponding ML pipeline stage.
@@
To ensure reliability in our categorization, the first two authors conducted an initial pilot study on an initial set of 65 samples to independently provide labels, and later discussed and resolved any classification disputes until consensus was reached. The inter-rater agreement, measured by Cohen’s Kappa, was 81% among the first two authors for a subset of 65 comments. This level of agreement, surpassing 70%, is considered adequate (byrt1996good). Following this pilot phase, we refined our categorizations by 1) eliminating any false positives, 2) removing the SATD instances labeled “unknown” and 3) removing the SATD from Non-ML components. Eventually, we found a total of 370 samples corresponding to different stages of the ML pipeline.
@@ 4.5.4. RQ4: How long does SATD survive in ML-based systems?
We perform survival analysis (miller2011survival) to examine, for each project, (i) the time until the introduction of SATD and (ii) the time taken to remove SATDs after they are introduced in a file for the first time. The former analysis examines the survival of ML software projects concerning debt occurrence, while the latter analysis is about the survival of SATDs instances once they are introduced in an ML-based software project.
Survival analysis (miller2011survival) is a statistical technique for analyzing the expected duration of time until the occurrence of an event of interest or censoring of an event (if the event does not happen). The event of interest in our study is the occurrence of SATD across the entire project history, and the length of the observation window is the start of the project until the time at which we obtained the projects’ latest snapshot at the time of performing the empirical study (Summer 2021). If the event is not observed during the observation period, the corresponding subject will be censored at the end of the period, which survival analysis is able to deal with.
@@
Time to event and status are two important variables for survival analysis. Time to event ($T$) is the elapsed time between the beginning of the observation (start date of the project) and the occurrence of the event of interest or the censoring of data. In our study, this time is measured in #days. Hence, $T$ is a random variable with positive values (miller2011survival). A summary of the Time ($T$), Event of interest, and Status is provided in Table 4. For the introduction of SATD, $T$ is the elapsed time between the start date of the project to the time of contamination of the file by SATD. For the removal of SATD, $T$ is the elapsed time between the removal and the introduction of SATD.
@@
The Survival function S(t) gives the probability ($P(T>t)$) that a subject will survive beyond time $t$. After arranging our data in increasing order of $T$, we plot the survival curve and estimate the survival probability using the non-parametric Kaplan-Meier estimator (kaplan1958nonparametric). The Kaplan-Meier estimation is computed following Equation 1, where $t_{i}$ is the time duration up to event-occurrence point $i$, $d_{i}$ is the number of event occurrences up to $t_{i}$, and $n_{i}$ is the number of subjects that survive just before $t_{i}$. $n_{i}$ and $d_{i}$ are obtained from the aforementioned ordered data.
@@ 4.5.5. RQ5: What are the characteristics of long-lasting SATDs in ML-based systems?
Our dimensions ‘Change’, ‘Diffusion’, ‘Complexity’, and ‘History’, are widely employed in the domain of SDP (software defect prediction), a well researched field in Software Engineering. The rationale for their usage in the context of explaining SATD survival is based on the understanding that areas of code that undergo frequent changes, or are complex, are more prone to defects. Given that SATD indicates sub-optimal implementations that may not be outright defects but could result in compromised software quality, we consider using these metrics for SATD prediction. Moreover, prior research in SATD prediction by Yan et al. (yan2018automating) previously employed such dimensions (i.e., diffusion, history, and message) in their study. In contrast to Yan et al.’s research, we exclude the message dimension from our study as our focus is on assessing the attributes of persistent self-admitted technical debt (SATD) rather than SATD detection itself.
@@
Assigning Class Labels.
To identify long-lasting SATDs, we compute the time elapsed since the creation of the comment admitting the debt. We compute this elapsed time using information from the commit diffs (i.e., ++++ and -- lines of code comments for each source code file in the commit). Specifically, we calculate the time between the introduction of the SATD comment and its removal. From our collected data, we observed cases in which the same SATD commit is present multiple times within the same file, and removed these cases from our case study to avoid biasing our results with duplicated information. After computing SATD duration time for a total of 6,427 SATD instances in our dataset, we computed the quartiles of the obtained distribution and used the first and fourth quartile to identify “quick removal” and “long lasting” SATD respectively. In total, we obtained 2,595 SATD instances with duration times above 75th percentile. We assigned them to the group of long-lasting SATDs. We obtained the first quantile, i.e., 2,595 instances of with duration times below the 25th percentile as the quick removal SATD.
Correlation and Redundancy Analysis.
@@
Model Validation.
To ensure the statistical robustness of our results, we perform a 100-out-of-sample bootstrapping. The approach involves sampling the training set from the whole dataset with replacement. The testing set comprises all the samples present in the original data but absent in the bootstrapped training data. We build and test models on each training and testing set to draw statistically robust conclusions as suggested by (tantithamthavorn2016empirical)
Model Interpretation.
We use SHAP (SHapely Additive exPlanations), to interpret and explain our models. A Shapley value is the average marginal contribution of a feature value across all possible coalitions (shapley201617). We compute these values using the Python framework, SHAP [^16] Accessed on: 2024-03-22, https://github.com/slundberg/shap which provides a collaborative game theory approach for the calculation of the average marginal contribution of each of the model’s features.
@@ 4.6. Replication Package
All the scripts along with the mined data is provided in the replication package [^17] Accessed on: 2024-03-22, https://drive.google.com/drive/folders/1n-gwAxFANS-PPewnk1Qgmnwd0zMAr7Yb *This tentative g-drive link will be replaced by a github repository after the reviewing process. .
@@ 5.1. RQ1: What is the prevalence of SATD in ML-based systems?
Figure 5. Percentage of SATD in ML and non-ML projects. The comparison has been made by collecting 424,248 bootstrapped samples from ML and non-ML comments and repeating the process 1,000 times (to ensure robustness).
@@
Machine learning-based projects have 2 times more SATD in comparison to non machine learning projects. Figure 5 shows the distribution of SATD in ML and non-ML software. The median of percentages of SATD in ML and non-ML software is 1.87% and 0.92%, calculated over 1000 times using bootstrapping, i.e., sampling with replacement of 424,248 (size of the ML-dataset) comments from both ML and non-ML categories. This indicates a 2 times more debt in ML projects. The comparisons provided in Figure 5 are statistically significant with the Wilcoxon Signed Rank test $p$-value lower than 0.05 and a large Cohen’s $d$ effect size difference. When considering all the comments in ML and non-ML code from our dataset (without any sampling), 1.87% of comment declarations in ML projects contain a technical debt admission (i.e., SATD), compared to only 0.93% for non-ML projects, having a ratio of 2.02. Our findings for SATD in non-ML software are aligned with the findings of prior research. For instance, Bavota and Russo (bavota2016large) found SATD in 0.3% of the comments of non-ML Java-based projects, which is comparable to the 0.9% obtained by our analysis for Python-based non-ML projects.
Summary of RQ1 Machine learning projects have a median percentage of SATD that is twice the median percentage of SATD in non-machine learning projects.
5.2. Discussion: Explanatory analysis for RQ1